OpenStudy (anonymous):

factor the following terms using a factor tree 1. 22x^2y 2. 35a^3b^2

OpenStudy (anonymous):

22x^2y=2*11*x*x*y 35a^2b^2=5*7*a*a*b*b

OpenStudy (yuki):

basically it's asking you to do prime factorization